### v1

# Learning to Debug: LLM-Organized Knowledge Trees for Solving RTL Assertion Failures

Abstract
Debugging is the dominant cost in modern hardware verification, where assertion failures are among the most frequent and expensive to resolve. While Large Language Models (LLMs) show promise, they often fail to capture the precise, reusable expertise that engineers apply, leading to inaccurate responses. We propose Grove, a hierarchical knowledge management framework that learns and organizes reusable debugging expertise into an LLM-organized knowledge tree for solving assertion failures. Grove distills debugging knowledge from prior cases and organizes it into a vertical tree of configurable depth, with each node encoding a concise knowledge item and explicit applicability conditions. During training, Grove uses a parallel, gradient-free loop where an LLM proposes tree modifications as structured JSON edits by learning from the cases. At test time, a budget-aware iterative zoom is performed to navigate the tree, retrieving a small set of applicable knowledge items that guide a base LLM’s hypothesis generation and fix proposals. Evaluated on a suite of assertion-failure cases, Grove delivers consistent gains in pass@1 and pass@5, demonstrating the value of structured knowledge evolution.
I Introduction
Modern hardware design relies on extensive verification to ensure correctness before fabrication [foster2020wilson]. Verification at scale produces many failures, yet debugging—isolating root causes and crafting RTL fixes—remains the dominant cost. Among failure modes, assertion failures are especially costly [zhou2025insights], requiring engineers to interpret logs/waveforms, reason about timing/causality, and implement precise RTL edits. LLMs and agents are increasingly used in EDA [thakur2023autochip, wei2025vericoder], but naïve prompting or flat retrieval rarely reproduces the domain-specific, reusable debugging expertise that practitioners rely on.
This expertise is concrete and action-oriented, targeting patterns that generic models miss, such as correcting spec-to-RTL constant mismatches, fixing off-by-one address predicates, repairing stateful counters, restoring XOR feedback, and zero-extending byte-to-word reads. In prior studies [zhou2025insights, bai2025fvdebug], LLM/agent systems often mislocalize faults, propose unstable or over-fitted edits, and bloat context with irrelevant snippets—leading to inaccurate responses and failed debugging. Absent a way to learn, structure, and validate such reusable knowledge from past cases, performance remains brittle and inefficient.
Figure 1: Overview of Grove: A knowledge management framework that transforms raw debugging experiences into reusable expertise. Training cases (left) undergo golden-aware reflection to distill validated knowledge items, integrated via atomic JSON edits into an LLM-organized hierarchical knowledge tree (center). Test cases (right) navigate this tree through budgeted snapshot-and-zoom operations to retrieve applicable knowledge, augmenting prompts with targeted debugging guidance. The tree maintains vertical organization (general to specific) and horizontal discipline (via shape guards) to scale across a large amount of training cases while preserving fast, precise retrieval.
This presents two fundamental challenges. First, knowledge acquisition: How can this latent debugging expertise be automatically learned from unstructured sources like past bug fixes, failure logs, and engineers’ experience, moving beyond ad-hoc case matching? Second, knowledge organization: How can this learned knowledge be organized for efficient retrieval and reuse? A flat, unstructured store is demonstrably suboptimal: it pollutes the LLM’s context with noisy information, scales poorly, and fails to match the hierarchical way engineers reason (e.g., decomposing an abstract pattern into concrete strategies and fixes).
To solve this, we present Grove (Governed Retrieval Of Validated Expertise), a hierarchical knowledge management framework for solving assertion failures in RTL designs. Grove distills reusable debugging knowledge from prior cases and organizes it into an LLM-organized knowledge tree. Instead of a fixed hierarchy, Grove dynamically structures knowledge into a tree of configurable depth. Each node encodes a concise actionable knowledge item, and explicit applicability conditions (apply_conditions) to govern its use.
During training, Grove runs a parallel, gradient-free learning loop. Each training case produces one or more candidate knowledge items, which are validated by injecting each item into the LLM prompt, generating RTL fixes, and checking them against the assertion using a model-checking tool. Only beneficial, performance-preserving knowledge is committed via LLM-generated JSON edit scripts—the LLM autonomously decides where to insert, update, move, or deprecate nodes, thereby organizing and refining the hierarchical tree under structural constraints. At test time, Grove navigates the tree using an LLM-guided iterative zoom on a budgeted snapshot of the tree, filtering candidates by applicability before assembly into a safe, case-specific knowledge set.
Scope. We instantiate and evaluate Grove on assertion-failure debugging in RTL. Extension to other debug tasks is conceptually supported by the framework but out of scope for this paper.
Our contributions can be summarized as follows:
  • An LLM-organized hierarchical knowledge tree with explicit applicability constraints and strict JSON edit operations. To the best of our knowledge, Grove is the first to combine per-item, tool-validated knowledge acquisition with a governed hierarchical memory tailored to RTL assertion-failure debugging.
  • A novel LLM-guided zoom procedure for test-time retrieval and a gradient-free, parallel learning workflow for scalable acquisition and organization of hierarchical debugging knowledge.
  • Comprehensive evaluation on SVA-Eval, demonstrating +12.0% absolute improvement in pass@1 on LLaMA-3 and +6.6% on o3-mini, with consistent pass@5 gains and ablations validating design choicea.
II Related Work
II-A Hardware Debugging and LLMs for EDA
Hardware debugging has evolved from manual inspection to sophisticated automated approaches. Formal verification [kern1999formal] and assertion-based verification [witharana2022survey] established foundational methods, with a recent survey [wu2024survey] analyzing ML integration across the verification spectrum. ML-driven bug localization [stracquadanio2024veribug, wu2024knowledge] achieves improvements in identifying faulty regions. Automated repair tools [ahmad2022cirfix, laeufer2024rtl] address debugging bottlenecks through genetic programming, symbolic execution, and LLM-based approaches. Hardware fuzzing techniques [laeufer2018rfuzz, jayasena2025fuss] systematically explore state spaces for corner-case bugs. Recent work integrates LLMs into EDA workflows [ho2025verilogcoder, bai2025fvdebug], though domain adaptation remains critical.
II-B Learning and Accumulating Knowledge
Valuable expertise for assertion debugging is latent in fixes, logs, and reviews. Beyond static retrieval, self-learning methods distill knowledge from feedback via reflection [shinn2023reflexion, madaan2023self] or persistent, evolving memory stores [madaan2022memory, zhong2024memorybank, zhang2025survey]. These stores can be dynamic graphs [xu2025mem], executable skill libraries [wang2023voyager], or architectural augmentations [wang2023augmenting, langgraph2024memory]. Grove differs by focusing on governed acquisition for hardware debugging: knowledge is mined from cases, validated for non-harmfulness, and committed via atomic edits with applicability contracts, building a domain-specific, dynamic tree rather than an unstructured general memory.
II-C Organizing and Retrieving Knowledge
RAG grounds LLMs [lewis2020retrieval], but flat case retrieval, common in EDA [qi2025verirag, wang2025veridebug, qayyum2025llm], struggles with scale and alignment to hierarchical reasoning. Structure-aware retrieval addresses this using document trees [zhang2024hierarchical, jiao2025hirag, huang2025retrieval] or knowledge graphs [edge2024local, jiang2025ras]. This aligns with broader AI trends of deliberate search [yao2023tree, hogan2021knowledge] and neuro-symbolic composition [karpas2022mrkl]. Related data system work targets fixed schemas [freire2025large, lao2025gptuner] rather than evolving domain knowledge. In contrast, Grove learns its own dynamic, LLM-organized knowledge tree, combining RAG’s grounding with agent memory’s adaptivity.
III Methodology
III-A Task Formulation
We formulate assertion-failure debugging as a two-regime learning and inference task. Figure 1 outlines the overall Grove framework.
Training cases
Each training sample consists of (i) a structured problem context and (ii) a golden solution. Following the methodology of prior work [zhou2025insights], the context includes the buggy RTL code, SystemVerilog assertion specifications, and assertion failure logs; the golden solution provides the buggy line(s) and corresponding fix.
During training, Grove exposes the golden solution only for reflection and validation; the LLM is asked to distill reusable debugging knowledge—concise, actionable statements paired with explicit apply_conditions that govern when they apply in natural language (detailed in Section III-D).
Test cases
Each test sample provides the same structured context without a golden solution. Grove retrieves a small set of relevant knowledge items using the budgeted snapshot+zoom protocol described in Section III-C (see Algorithm 1), and appends the retrieved knowledge to the test prompt for a base LLM to produce hypotheses and fixes. This zoom-then-augment design yields improved robustness by grounding the LLM’s reasoning in validated, case-relevant debugging expertise.
III-B Overall Framework
Grove addresses the dual challenges of knowledge acquisition and knowledge organization by learning reusable, validated debugging knowledge from past cases and maintaining it as a vertical, LLM-organized knowledge tree. Each node (knowledge item) carries (i) a concise, actionable statement, and (ii) explicit apply_conditions that govern when the statement should be applied. The apply_conditions are written by the LLM in natural language following a schema that asks for concrete syntactic and semantic cues (e.g., specific operators, module names, signal widths).
Abstraction is emergent: upper levels tend to encode more general regularities, while deeper levels capture increasingly specific realizations; this is regulated by these applicability constraints and edit-time validation. As the tree grows, Grove enforces simple structural constraints on depth and branching to keep it navigable and efficient; we describe these safeguards in more detail in Sections III-D and III-F.
III-C Handling Growing Tree Size: Snapshot and Zoom
Motivation. A successful deployment of Grove will accumulate hundreds or thousands of nodes. Exposing the entire tree at once would exceed context budgets and drown useful knowledge in noise. We therefore separate what is stored (the full tree) from what is shown (a budgeted view), and give the LLM a small set of structured zoom operations to access local detail on demand.
Budgeted snapshot. Grove constructs a budgeted structural snapshot of the knowledge tree. This snapshot preserves the global shape of the tree while respecting a token budget $B_{\text{snap}}$, and is created by including a representative subset of nodes from each level. For each included node, we render its identifier, a compact summary, and its apply_conditions. Nodes with unexpanded children are marked with ellipsis indicators (e.g., “…”), signaling to the LLM that zoom operations can reveal additional detail. This design balances breadth (global tree shape) with depth (local detail on demand).
JSON-based interaction protocol. All LLM–tree interactions use a structured JSON protocol. On each round, Grove synthesizes a prompt containing the current context $C$, the snapshot $S$, any previously expanded subtrees $V_{\text{details}}$, and a JSON response schema.
The LLM responds with a JSON object that may contain two types of fields: (i) read operations—a read_ops list specifying zoom actions such as expand_node (for subtree expansion) or list_children (for immediate descendants)—and (ii) candidate selections—a select_node_ids list marking promising knowledge items. Read operations are executed to generate detailed textual views, which are accumulated across rounds under a per-round chunk budget $B_{\text{chunk}}$.
The semantics of select_node_ids are task-specific: at test time, these are the node IDs for retrieval. During training, Grove uses an analogous JSON protocol, but instead of select_node_ids the LLM returns a JSON edit script specifying tree modifications (insert_node, update_node, move_node, deprecate_node) with explicit node references and field updates.
Protocol workflow. Algorithm 1 formalizes the retrieval protocol, which follows the iterative expansion concept depicted in Figure 1. The process begins by providing the LLM with an initial, high-level view of the tree: the budgeted snapshot $S$ (Line 1), which acts as a navigable map where complex branches are collapsed. The system then enters a loop for up to $R$ rounds.
In each round, the LLM examines its current view of the tree—the snapshot plus any detailed views requested so far—and can issue read_ops to “zoom in” on specific nodes. When the LLM requests an expansion, the system generates a more detailed textual representation of that subtree and adds it to the set of detailed views. On the next iteration, the LLM is presented with an updated view that combines the global snapshot with these expanded regions, allowing it to progressively move from a general overview to specific details.
The loop terminates once the LLM stops requesting further zooms (or the round budget is reached). At that point, the system aggregates all proposed select_node_ids, and returns a final set of node IDs $\hat{\mathcal{K}}$ (Line 15). During training, the same iterative exploration logic applies, but the final JSON object contains an edit script instead of node IDs.
Algorithm 1 Grove Snapshot+Zoom Retrieval Protocol
0: Context $C$, tree $T$, snapshot budget $B_{\text{snap}}$, max rounds $R$
[…]
2:for $r=1\ldots R$ do
3:  $\text{p}\leftarrow\textsc{ConstructPrompt}(C,S,V_{\text{details}})$
4:  $(\texttt{read\_ops},\texttt{select\_node\_ids})\leftarrow\textsc{LLM}(\text{p})$
5:  if select_node_ids is non-empty then
6:   $C_{\text{cand}}\leftarrow C_{\text{cand}}\cup\texttt{select\_node\_ids}$
7:  end if
8:  if read_ops is non-empty then
9:   $V_{\text{details}}\leftarrow V_{\text{details}}\cup\textsc{ApplyRead}(T,\texttt{read\_ops})$
10:  else
11:   break
12:  end if
13:end for
[…]
15:return $\hat{\mathcal{K}}$ {Final set of candidate node IDs}
III-D Training-Time Knowledge Acquisition and Validation
Parallel learning workflow. To scale knowledge acquisition, Grove leverages a gradient-free, parallel learning architecture. The system spawns multiple concurrent workers (via multi-processing), where each worker processes a disjoint subset of training cases. All workers share a single underlying knowledge tree data structure.
Golden-aware reflection and tree edits. For each training case, Grove performs a golden-aware reflection that inspects the case context alongside the golden fix. Using the snapshot+zoom protocol, the LLM first explores the existing tree to locate relevant regions and potential parents. The reflection step concludes with the LLM proposing a candidate tree edit, which it formulates as a JSON edit script as described in Section III-C.
Each script lists one or more operations—insert_node (add new knowledge), update_node (modify existing content or metadata), move_node (restructure hierarchy), or deprecate_node (mark obsolete)—along with explicit node references (by ID or path) and field updates. This structured, machine-checkable output forces the LLM to formalize its reasoning into executable operations and allows each worker in Grove to modify a shared tree data structure.
Per-knowledge-item validation. Before modification, the worker evaluates each candidate knowledge item individually extracted from the edit scripts. It regenerates solutions for the training case with that single item’s statement injected into the prompt. If the item does not degrade pass@k on its originating case, it is marked as non-degrading and becomes eligible for integration. This criterion is chosen over strictly enforcing improvement to enable knowledge generalization and to avoid discarding valid, widely applicable knowledge. Our current validation scheme uses the originating case, which keeps cost manageable. Extending validation to a separate held-out subset is an interesting future effort.
Governed integration. Validated scripts are parsed into discrete graph operations and applied atomically to the shared tree under a global write lock. This process is termed governed integration because Grove actively enforces structural invariants during execution: schema checks ensure all nodes possess explicit apply_conditions, and a verticality constraint forces nodes at level $l$ to attach only under parents at level (l1)(l{-}1). These checks ensure that abstraction levels progress cleanly from general (upper levels) to specific (lower levels).
III-E Test-Time Knowledge Retrieval
Embedding-based similarity alone is insufficient for RTL debug retrieval. It ignores applicability structure: whether a rule should apply under specific control, temporal, or module contexts. Prior work shows that naïve dense retrieval can pollute context and misalign with downstream tasks when applicability predicates matter [liu2024lost], and that tree-organized retrieval can better integrate information across abstraction levels [sarthi2024raptor].
In contrast, Grove leverages the LLM’s reasoning to navigate a budgeted snapshot and issue targeted zoom operations (JSON read_ops), guided by the learned apply_conditions. This lets the model connect hints present in the test context to rules whose applicability matches, rather than relying on surface-level textual similarity.
At inference, Grove follows the iterative snapshot+zoom protocol detailed in Algorithm 1. The LLM progressively accumulates detailed views via read_ops and marks candidates via select_node_ids. The retrieved items are appended as concise statements; apply_conditions guide selection and are surfaced in zoom views.
III-F Maintaining Shape and Scalability
To ensure the knowledge tree remains efficient and navigable as it grows, Grove uses active governance over its shape and records detailed operational telemetry. Scalability is primarily enforced by shape guards, which are strict, configurable checks applied during every write operation. Key parameters include budgets for the maximum number of root-level (level-1) nodes and the maximum children per parent node. By enforcing these constraints at edit time and asking the LLM to re-propose edits when the constraints are violated, Grove prevents uncontrolled branching and mitigates degradation in retrieval latency and quality as more knowledge items are added.
III-G Complexity and Resource Use
Let $N$ be the number of training cases, $M$ the average number of candidate knowledge items generated per case via golden-aware reflection, $E$ the cost of a single pass@k evaluation (one complete generation and evaluation pass), and $R$ the number of zoom rounds. Training cost is dominated by per-statement validation at $O(M\!\cdot\!E)$ per case; edit integration is near-constant per accepted item. Retrieval uses at most $R$ compact prompts, with local expansions touching only targeted subtrees. Parallel workers processing the training data in a multi-processing way yield near-linear speedups until external rate limits or prover/simulator throughput saturate.
IV Experiments
We evaluate Grove on the SVA-Eval benchmark of assertion-failure cases [zhou2025insights]. Code with data will be released for easy reproduction of our results.
IV-A Dataset
We adopt SVA-Eval [zhou2025insights], which provides specification (Spec), buggy code, logs, and a golden fix for each case. We combine the human-curated (SVA-Eval-Human) and LLM-generated (SVA-Eval-Machine) splits, then partition the union into 80% training and 20% testing, ensuring no module appears in both splits to prevent data leakage. SVA-Eval is produced via tool-validated augmentation: SVAs and bugs are generated and filtered with HDL compilers and formal/symbolic checks to ensure the bug–SVA pair triggers an assertion and the golden fix resolves it [zhou2025insights].
IV-B Baselines
TABLE I: Main evaluation results on the SVA-Eval test set. We report pass@k for all methods across two base LLMs for testing.
Method LLaMA-3 o3-mini
Pass@1 Pass@5 Pass@1 Pass@5
Zero-shot 0.727 0.754 0.847 0.851
Case-RAG 0.680 0.680 0.787 0.820
Ref-RAG 0.508 0.536 0.689 0.700
GlobalMem 0.732 0.765 0.836 0.847
Ref-seqR 0.749 0.781 0.880 0.885
Grove (No-Apply) 0.792 0.820 0.901 0.901
Grove (No-Val) 0.814 0.831 0.907 0.907
Grove (No-LLM-Ret) 0.754 0.781 0.885 0.891
Grove (No-Zoom) 0.803 0.831 0.880 0.885
Grove (Ours) 0.847 0.869 0.913 0.913
We compare Grove against knowledge management baselines spanning embedding-based retrieval, flat-memory, and LLM-guided selection: Case-RAG (Case-level RAG). Each training case is treated as a retrievable document. At test time, we retrieve the top-$k$ most similar cases using dense embeddings and include them in the prompt context. Ref-RAG (Flat knowledge RAG). Reflexion [shinn2023reflexion] distills debugging knowledge from iterative self-reflection into a flat knowledge store. Retrieval follows Case-RAG but operates over learned knowledge items rather than full cases. GlobalMem (Global memory). All Reflexion knowledge items are appended to every prompt without retrieval, testing whether unfiltered knowledge access suffices. Ref-seqR (LLM-based retrieval). Inspired by LLM-as-retriever approaches [asai2024self], the system chunks the Reflexion knowledge store and sequentially prompts the LLM to select relevant items, without Grove’s hierarchical organization. For Case-RAG and Ref-RAG, we vary $k\in\{1,2,4,8\}$ and report the best result in Table I for conciseness.
IV-C Implementation Details for Grove
For Grove, the knowledge tree is constructed following the workflow described in Section III. The snapshot+zoom protocol is governed by two main budgets: the initial snapshot budget is set to $B_{\text{snap}}=80000$ tokens, and the per-round chunk budget is $B_{\text{chunk}}=12000$ tokens. During testing, we execute at most $R=10$ zoom rounds. The tree’s structure is managed by shape guards that cap the number of root-level nodes at 216 and the fan-out for any parent node at 144.
IV-D Evaluation Protocol
To evaluate the correctness of the generated fixes, we use Cadence Jasper (version 2023.12) as the model-checking tool. We apply the patch proposed by each model to the buggy RTL and run the verifier. A fix is considered successful if the modified code compiles and all assertions pass. Consistent with the SVA-Eval benchmark’s methodology [zhou2025insights], we report the pass@1 and pass@5 metrics. Testing is based on two base LLMs: Llama 3.3 70B Instruct [dubey2024llama] and OpenAI o3-mini [openai2025o3mini].
IV-E Results and Analysis
Table I presents the primary results on SVA-Eval. Grove delivers the top pass@k across both base LLMs, outperforming zero-shot LLaMA-3 and all non-hierarchical memory baselines. Case-RAG and Ref-RAG provide gains over zero-shot in some settings, confirming that retrieval can help; however, both lag behind Grove, indicating that naïvely surfacing entire cases or flat memories remains suboptimal. GlobalMem and Ref-seqR show that simply providing a strong global rulebook or using a flat LLM-based selector is not sufficient. Grove’s learned, hierarchical knowledge with explicit applicability conditions enables more precise, targeted retrieval.
IV-F Ablation Study
As shown in Table I, removing LLM-guided retrieval (No-LLM-Ret)—i.e., replacing snapshot+zoom with embedding-based similarity search—causes the largest degradation in pass@1, indicating that semantic similarity alone cannot capture the nuanced applicability conditions needed for effective selection.
Removing applicability conditions (No-Apply) from the tree also harms performance, confirming that explicit, natural-language predicates are essential for precise routing; and dropping the iterative zoom protocol (No-Zoom) underperforms the budgeted, progressive exploration used by Grove.
Finally, removing per-statement validation (No-Val) further reduces accuracy. Together, these ablations confirm that Grove’s gains arise from the synergy of validated knowledge, hierarchical organization, explicit applicability, and LLM-guided retrieval.
IV-G Retrieval Quality Study
Beyond ablations, we additionally evaluate retrieval quality with an LLM-as-judge [es2024ragas], which sees the case context and a method’s retrieved items and returns a calibrated helpfulness score $[0,1]$. We also judge whether a method’s final answer is relevant (AnsRel) and supported (AnsSup) by its retrieved items. Table II shows Grove retrieves more helpful items, and its answers are more often relevant and supported than non-hierarchical baselines.
TABLE II: Retrieval quality (LLM-as-judge) on LLaMA-3. Higher is better.
Method Helpfulness AnsRel AnsSup
Case-RAG 0.221 0.419 0.193
Ref-RAG 0.315 0.471 0.281
GlobalMem 0.279 0.440 0.215
Ref-seqR 0.441 0.587 0.368
Grove 0.626 0.711 0.539
IV-H Characteristics of the Learned Knowledge Tree
We analyze Grove’s learning process and the resulting tree structure. Our gradient-free, parallel workflow is efficient: with 8 parallel workers, the full knowledge tree was learned from the 80% training split (732 cases) in under 24 hours.
Figure 2 visualizes the tree’s growth dynamics. We observe steady, monotonic accumulation at every level, indicating that Grove continuously distills and integrates new expertise rather than oscillating or collapsing categories. The right panel provides a structural view of the final tree, illustrating the hierarchical organization across six levels. The majority of knowledge resides at L2 and L3, representing core knowledge.
The emergent hierarchy—organized by the LLM rather than hand-designed—arises from governed learning with per-statement validation, explicit applicability conditions, and shape guards that maintain depth and fan-out discipline.
Figure 2: Knowledge tree growth and structure. Left: Stacked area of nodes per level across training steps; the upper envelope gives the cumulative node count, demonstrating steady, continuous learning. Right: Final tree layout showing hierarchical organization across six levels.
IV-I Data Efficiency and Computational Cost
To study robustness under limited training data, we vary the fraction of SVA-Eval training cases used to train Grove, keeping the test set fixed (Figure 3). Both pass@1 and pass@5 improve steadily as more data are available, indicating that Grove remains effective under data scarcity and continues to benefit from additional cases. Across all methods, end-to-end latency per case including inference and evaluation is dominated by Jasper runs; the additional LLM calls introduced by Grove keep total time within the same order of magnitude as the strongest baselines.
Figure 3: Effect of training set size for Grove with LLaMA-3.
IV-J Case Study: Training-Time Knowledge Acquisition
Figure 4 illustrates Grove’s knowledge acquisition process. Given a training case where the register width mismatches the specification, the LLM performs golden-aware reflection and proposes a JSON edit script to insert a new level-2 knowledge item under an existing parent. The proposed item then undergoes per-statement validation (pass@k unchanged on the originating case) before atomic integration via the structured edit operation.
Training Case: dataGenerator
Assertion
@(posedge clock) disable iff(!nReset)
!testModeFlag |=> (adcData == \$past(adc_databus));
Spec (excerpt)
## 2. Module Interface
### Inputs
- **adc_databus** (10-bit): Input bus carrying data from an ADC.
Buggy RTL (excerpt)
reg [8:0] adcData;
assign dataOut = testModeFlag ? testData : adcData;
always @(posedge clock, …
Golden fix
reg [9:0] adcData;
(a) Training case with golden solution.
LLM-proposed JSON edit (excerpt)
"ops": [{
”type”: ”insert_node”,
”parent_ref”: {”id”: …},
”node”: {
”level”: 2,
”title”: ”DataGenerator Width Mismatch in ADC Data Capture”,
”knowledge_statement”: ”Ensure that capture registers use the exact width specified for their input buses to prevent data capture assertion failures.”,
”apply_conditions”: ”Applicable when an ADC capture register is declared narrower than the specified bus width (e.g., reg [8:0] for a 10-bit bus). Look for width mismatches between RTL registers and spec/properties.”
(b) Proposed edit and validation.
Figure 4: Training-time knowledge acquisition. (a) Golden-aware reflection on dataGenerator. (b) LLM proposes inserting a level-2 knowledge item.
IV-K Case Study: Test-Time Knowledge Retrieval
Figure 5 presents a case study on an Avalon-interface FIFO counter with memory-mapped address decoding. The bug is subtle: the RTL uses bitwise AND (&) instead of a zero-extending operation, causing the readdata to be all zeros instead of a zero-extended value. As shown, Grove-LLaMA-3 retrieves a highly relevant knowledge item about zero-extension. Guided by this, it produces a correct fix, while the baseline, lacking this context, misdiagnoses the problem.
Parent knowledge item (L1). Ensure that both read ports in register-file modules conditionally output zero for address 0.
apply_conditions: This rule now serves as a grouping node for interface assertions related to register file and memory modules …
Children (L2):
    Ensure any write operation sets register 0 to zero …
    Retrieved knowledge item ($\bigstar$).
    “Read multiplexer assignments must use (address == 0) to route valid data, ensuring that data_out is zero-extended for non-zero addresses per specification.”
      apply_conditions: When constructing the read multiplexer, check the address condition so that the produced data matches …
    Replace the incorrect bit slice assignment [4:1] with [3:0] to match the assertion requirement …
(a) Knowledge tree context around the retrieved knowledge item ($\bigstar$).
Assertion
address == 0 |=> readdata[5:0] == \$past(in_port)
&& readdata[31:6] == 0;
Buggy RTL (line 32)
readdata <= {32’b0 & read_mux_out};
Bitwise AND masks read_mux_out to 0
Golden fix (from dataset)
readdata <= {32’b0 | read_mux_out};
Bitwise OR with 32’b0 zero-extends the 6-bit value to 32 bits
Baseline fix (zero-shot)
assign read_mux_out = (address == 0) ? data_in : 6’b0; …
Jasper: FAILED (wrong location)
Grove fix (with knowledge)
readdata <= {26’b0, read_mux_out};
Jasper: PASSED. In this module, both this fix and the golden form a 32-bit zero-extension of the 6-bit read_mux_out.
(b) Fix comparison.
Figure 5: Test-time knowledge retrieval. (a) Learned knowledge visualization. (b) Effect on fix generation.
V Conclusion and Future Work
We presented Grove, a hierarchical knowledge management framework for RTL assertion-failure debugging. Our work demonstrates that an LLM can effectively maintain and organize a dynamic, hierarchical knowledge base. This gradient-free learning process allows the knowledge tree to be organized by the LLM as a structured, symbolic store, integrating validated expertise via atomic edits rather than costly backpropagation. Experimental results demonstrate Grove’s effectiveness, showing pass@1 and pass@5 improvements over baselines.
Future work could extend Grove to other debugging tasks (e.g., value mismatches, timing violations) and broader EDA tasks (e.g., RTL generation, test synthesis, coverage closure). Key research challenges include: distilling robust knowledge in complex design scenarios, enabling continual learning for evolving codebases, and further reducing retrieval latency. Addressing these challenges would unlock more capable and efficient agents across design automation.